The following identity may be regarded as a formulation of the Weierstrass product for the Gamma function.

∫_{0}^{∞} t^{x}e^{−t} \frac{dt}{t} = \frac{1}{x} ∏_{k=1}^{∞} \frac{(1 + \frac{1}{k} )^{x}}{(1 + \frac{x}{k})}

Understanding the derivation of this identity is reasonable for a bright student of first year undergraduate calculus in the United States.